The sum of first 20 terms of the series 1,4,9,16.......
Answer by satyareddy22(84)

sum of squares of first n natural numbers = n(n+1)(2n+1)/6
n=20
= 20(20+1)(40+1)/6
=20(21)(41)/6
=2870